VIF En français | In english
Home | My account | Support | Contacts
Saturday, August 19 2017 18:15:45 EDT
search for on  
This page contains a summary of several connexion tests performed on key components of the network.

Server Status (HTTP) Good (SMTP) Good (DNS) Good (DNS) Good (POP3) Good (Webmail) Good (Domain reg.) Good
Extern. 1 Good
Google (HTTP) Good
Yahoo (HTTP) Good (IMAP) Good (FTP) Good
Extern. 2 Good
Hotmail (HTTP) Good
ICQ/AOL (5190) Good
Last update on Sat Aug 19 18:15:01 2017

Network status history:

2017-05-30 11:50:14 EDTAll systems are running smoothly
2017-05-14 19:00:00 EDTIn order to limit the propagation of the current SMBv1 exploits:
The folowing ports are blocked: UDP ports 137-138 and TCP ports 445,137,139
2017-03-10 10:57:48 ESTAll systems are running
2017-02-16 12:33:03 ESTAll our systems are functioning properly.
2017-01-19 15:04:47 ESTOur phone system rebooted. If you have issues restart your VoIP adapter. All systems are functioning properly.
2016-12-31 17:37:40 UTCOn this last Saturday of 2016 our servers are working fine and no outage was reported. The email service is running with no issue. VIF is wishing you a Happy New Year!
2015-08-29 17:56:50 UTCUPDATE: Secondary maintenance will be done with the power source in downtown center. Work must be done between 2:30pm and 3:00pm and our services will be turned off. Please make sure to restart all your devices after this time to insure you get back the service quickly
2015-08-29 08:44:55 UTCMaintenance work was done in downtown in our main colocation center and power was turned off for 6 hours. Services have been back since 6:30 am. If you are still experiencing loss of service, please restart all your equipments by turning off and turning on. Please be careful not to reset your devices
2010-04-11 06:00:00 EDTOn Sunday April 11, we experienced a major outage resulting from a general equipment failure. The service has been re-established since then but we are currently experiencing congestion issues, as many of our users are trying to reconnect at the same time. At this point in time, the best solution is to keep trying to reconnect until authentication succeeds. Thank you for understanding.
2009-11-16 10:43:13 ESTThe webmail service is now fully functionnal.

VIF Internet 368 Notre Dame O suite 200 Montreal QC H2Y 1T9 - 514-353-9988 / sans-frais: 888-321-3737